Part A What is the magnification of an astronomical telescope whose objective lens has a focal...

Part A

What is the magnification of an astronomical telescope whose objective lens has a focal length of 73 cm and whose eyepiece has a focal length of 2.6 cm ? Follow the sign conventions.

Part B

What is the overall length of the telescope when adjusted for a relaxed eye?
Answer #1

Part A

Magnification of telescope is given by,

m = -fo/fe

here, fo = focal legth of objective lens = 73 cm

fe = focal length of eyepiece lens = 2.6 cm

then, m = -73/2.6 = -28.08

In two significant figure:

m = -28

Part B

Length of the telescope will be:

L = fo + fe

L = 73 + 2.6

L = 75.6 cm

In two significant figure:

L = 76 cm

"Let me know if you have any query."